Please solve this word problems for me. I need the algebra equation along with the answer, please.
Mardi received an inheritance of $60,000. She invested part at 8% and deposited the remainder in tax-free bonds at 7%. Her total annual income from the investment was $4700. Find the amount invested at 8%.
thanks

Answer by Earlsdon(6294)

Let x = the amount Mardi invested at 8% and $60,000 - x = the amount invested in the tax-free bonds at 7%.
The interest received from the 8% investment is: 0.08x
The interest received from the tax-free bonds is 0.07($60,000 - x)
The total interest is:
0.08x + 0.07($60,000 - x) = $4,700 Simplify and solve for x.
0.08x + $4,200 - 0.07x = $4,700
0.01x + $4,200 = $4,700 Subtract $4,200 from both sides.
0.01x = $500 Divide both sides by 0.01
x = $50,000 The amount invested at 8%